Try evaluating the numerator and denominator separately.]
(c) Sketch a graph of cosh θ vs. θ, letting θ range from large negative
values to large positive values. [Hint: Note that for large positive θ,
e
−θ
is very small (e
−θ ≈ 0).]
(d) Sketch a graph of sinh θ vs. θ, letting θ range from large negative
values to large positive values. [Hint: Note that for large positive θ,
e
−θ
is very small (e
−θ ≈ 0).]
(e) Sketch a graph of v/c vs. θ for v/c = tanh θ, letting θ range from
large negative values to large positive values. [Hint: Note that for
large positive θ, e
−θ
is very small (e
−θ ≈ 0).]
(f) Let τ = 1sec and let θ range from large negative values to large
positive values. Sketch the curve in spacetime described by:
t = τ cosh θ,
x = (cτ) sinh θ.
(g) Let d = 1Ls and let θ range from large negative values to large
positive values. Sketch the curve in spacetime described by:
t = (d/c) sinh θ,
x = d cosh θ.
1
Expert's answer
2019-12-06T09:39:55-0500

c) The graph of the function is shown below for large negative theta to large positive theta:


"y(\theta)=\text{cosh}\theta"


d) Below is the graph for


"y(\theta)=\text{sinh}\theta:"


e) Below we can see the graph of

"y(\theta)=\text{tanh}\theta:"


f) The following are the graphs of the two functions:


"y(\theta)=\tau\text{cosh}\theta=\text{cosh}\theta\space\text{(red)},\\
g(\theta)=c\tau\space\text{sinh}\theta=3\cdot10^8\cdot\text{sinh}\theta\space \text{(green)}:"


g) Finally, show the graphs


"y(\theta)=\frac{d}{c}\text{sinh}\theta=\text{sinh}\theta\space\text{(red)},\\
g(\theta)=d\space\text{cosh}\theta=3\cdot10^8\cdot\text{cosh}\theta\space \text{(green)}:"